At a rock concert, a dB meter registered 135 dB when placed 2.3 m in front of a loudspeaker on the stage. The intensity of the reference level required to determine the sound level is 1.0 x 10-12 W/m2 


 Part A 

What was the power output of the speaker, assuming uniform spherical spreading of the sound and neglecting absorption in the air? Express your answer to two significant figures and include the appropriate units.


 Part B 

How far away would the sound level be 86 dB? 

Express your answer to two significant figures and include the appropriate units.
Answer #1

Part A

135 dB = 10*log(I/(1*10^-12))

10^(13.5) = I/(1*10^-12)

3.16*10^13 = I/(1*10^-12)

I = 31.6 W/m^2

Area = 4*pi*(2.3)^2 = 66.48 m^2

Power = Intensity*Area

= 31.6 W/m^2 * 66.48m^2

= 2100 W

Part B

86 dB = 10*log(I/(1*10^-12))

(10)^8.6 = I/(1*10^-12)

3.98*10^8 = I/(1*10^-12)

I = 3.98*10^-4 W/m^2

P = I*A

A = P/I

4*pi*r^2 = P/I

r = sqrt(2100/(3.98*10^-4*4*pi))

= 650 m